# Masked Sensory-Temporal Attention for Sensor Generalization in Quadruped Locomotion

### v1

@@ 5 Experiments and Results
We design and conduct various simulation experiments to evaluate the effectiveness of the proposed masking mechanism, overall performance of MSTA, and its generalization ability for different sensor data. We mainly focus on three metrics: linear velocity tracking return per step, angular velocity tracking return per step, and the total final reward return, which tells us how the agent can conduct the task (following users’ commands) and what the overall performance is. All reported results are averaged over 5000 trails with five terrain types and different levels. They are normalized on the basis of respect teacher data for easy comparison.
5.1 Impact of Mask Ratio
@@
First, we would like to investigate the maximum portion of missing data MSTA can handle to reconstruct robot states. During the transfer stages, we set the masking ratio to 0%, 25%, 50% and 75% independently and test the 16 model with these masking ratios. Figure 2 shows the resultant heatmap matrix. When trained without masking, despite the model having very good performance when all information is present, it suffers from missing data and cannot efficiently reconstruct the status. Also, we can see that the performance is more dependent on the masking ratio in the second stage than that in the first stage. This is because in the second transfer stage, the student is actually interacting with the environment to reduce the gap caused by missing information and observation shift. In contrast, the mission of the first stage is to generate a usable policy that outputs reasonable actions so the agent does not fail dramatically and has the chance in the second stage to generate high-quality trajectories for optimization, which is achievable even with a masking ratio of 75%. This also demonstrates the importance of using two-stage transfer.
@@
Comparing the performance of these models, we choose the one trained with the masking ratio of 75% in the first stage and 50% in the second stage, which can well balance the resource usage and performance. This selected model can restore almost full performance even under a masking ratio of 50% during testing. Such ratio can significantly reduce the training time and resources required.
@@
Figure 2: Heatmap matrix for the performance of models that are trained with different combinations of mask ratios. The three rows from top to bottom represent the linear velocity tracking, angular velocity tracking and total reward return respectively. The four columns are the masking ratios applied during testing. For each sub-figure, the y-axis is the masking ratio applied during the offline pretraining stage and the x-axis is the masking ratio applied during the online correction stage.
5.2 Comparison with Baselines
We compare MSTA with two baselines. The first is RMA [4], which is implemented with TCN [31] to capture temporal information. The second is TERT [16], a transformer-based framework with linear projection for tokenization of observations and actions in two favors: concatenated single token and separate tokens for states and action, resulting in $T$ and $2T$ tokens respectively [18]. Both RMA and TERT have access and need the full observation information. To further evaluate the importance and capability of the transformer structure in our method, we replace it with a GRU [32] model. For both MSTA and GRU, we apply a testing mask of up to 50%, as identified in Section 5.1. All baselines are trained with the same two-stage transfer method sharing a common well-trained teacher network. Table 1 shows the comparison results. When fully optimized with the teacher-student transfer, the performance of all trained policies with complete observations is very close, often within just 2% difference. When faced with incomplete information, transformer-based MSTA can have a better understanding of the data and reconstruct the robot state more accurately than the GRU-based network, even with only half of the information.
@@ 5.2 Comparison with Baselines
Table 1: Comparison results on different terrain types in terms of linear velocity tracking, angular velocity tracking and total reward return. For MSTA and GRU, we apply the mask ratios of 0%, 25% and 50% during testing to evaluate their robustness against incomplete information.
  Terrain Metric Ours RMA TERT GRU
0% 25% 50% Concat Seperate 0% 25% 50%
  Smooth Slope Linear Tracking 1.00 0.99 0.99 1.00 1.01 1.00 1.00 0.96 0.78
Angular Tracking 1.02 1.02 1.00 1.02 1.02 1.02 1.02 1.01 0.97
Total Reward 1.01 1.01 0.99 1.02 1.02 1.02 1.01 0.99 0.84
Rough Slope Linear Tracking 0.97 0.94 0.95 0.99 1.00 0.98 0.96 0.90 0.74
Angular Tracking 1.01 1.00 0.98 1.02 1.02 1.01 1.01 1.00 0.95
Total Reward 0.98 0.95 0.95 1.00 1.01 0.99 0.97 0.92 0.79
Stairs Up Linear Tracking 0.93 0.91 0.91 0.94 0.95 0.95 0.90 0.91 0.69
Angular Tracking 1.00 0.98 0.97 0.99 1.00 1.01 0.99 0.98 0.94
Total Reward 0.95 0.91 0.87 0.95 0.99 1.00 0.90 0.87 0.72
Stairs Down Linear Tracking 0.94 0.93 0.93 0.96 0.97 0.95 0.92 0.93 0.74
Angular Tracking 1.00 0.99 0.99 1.00 1.01 1.01 0.99 0.98 0.95
Total Reward 0.94 0.91 0.91 0.95 1.00 0.98 0.89 0.90 0.77
Discrete Linear Tracking 0.96 0.96 0.94 0.97 0.99 0.98 0.93 0.88 0.75
Angular Tracking 1.01 1.01 0.99 1.01 1.02 1.02 1.01 0.99 0.95
Total Reward 1.00 0.97 0.96 0.99 1.04 1.02 0.94 0.89 0.83
5.3 Generalization, Robustness and Flexibility
While achieving state-of-the-art performance, MSTA additionally offers the benefit of generalization and flexibility to customize the model after training or even on the fly to fit the deployment requirement. We can balance the performance and computation power required by randomly masking out the observation directly, just like we do during testing. As quadrupedal robots are often not equipped with high-power computers and need to be off the grid to work remotely, such balance is crucial. The in-context understanding capability of the transformer can give us more insight.
@@ 5.3 Generalization, Robustness and Flexibility
Figure 3: Performance after various setups: 3(a) certain sensory feedback is completely removed from the observation; 3(b) certain numbers of joint encoders are masked out; 3(c) different history time window $T$ is applied; 3(d) using minimized observations before and after finetuning; 3(e) using different numbers of self-attention blocks tested with a random mask of 50%.
@@
Important Sensory Feedback. As our random masking is applied to each dimension independently, even with a mask of 50%, the proprioceptive information is often not completely wiped out and the transformer can still trace it in the given time window. We further investigate the impact of removing one sensor completely from the feedback. The results are shown in Figure 3(a). It is clear that certain feedback like $\dot{q}$, $c$, $\omega$ and even $a_{t-1}$ are quite redundant and a well trained transformer-based MSTA can easily compensate the missing information.
@@
Missing of Sensor Dimension. For some proprioceptive information, it is provided by multiple sensors, such as the individual joint encoders and force senors on each foot, meaning that they can also be damaged independently due to wear and tear from daily operations and it is not easy to have a redundant sensor. Among these sensors, joint encoders are the source of both $q$ and $\dot{q}$ for the observation. From previous analysis, missing of joint information can be crucial. We investigate the scenario where only a few encoders are dead or the data are compromised and need to be excluded. We conduct the test to mask certain numbers of joint encoders. For each masked joint, the related $q$ and $\dot{q}$ are removed completely from the observation. The results are shown in Figure 3(b). The loss of the joint encoder can have a great impact on the performance as the related information is very essential for quadruped locomotion. However, our transformer model can still handle multiple missing encoders before large performance degradation.
@@
Time Window. Another configuration we can easily adjust on the fly for balance is the time window for historical information. The default window, $T=15$, is equivalently to past 0.3s. We check whether such a long sequence of information is necessary by applying different time windows without masking. The results are shown in Figure 3(c). It is clear that MSTA can efficiently extract and reconstruct the robot state for actions with only 7 steps of past information. Interestingly, given a longer timeframe like $T=20$ and the position the transformer has never seen during training, MSTA is still robust and not affected by these unknown information.
Unseen Information. We further assess the model’s capability to handle previously unseen information by appending the sequence with random dummy tokens. With the increasing of unseen information, it become more challenging for the policy in following commands. However, even with 256 dummy tokens, equivalent to a camera frame with ViT [30], the agent can still produce reasonable commands for exploration. This demonstrates that the model can be used as a solid foundation for further extension with high-dimensional information.
@@
Minimized Observation and Fine-tuning. From the previous analysis in this section, we have identify the important sensors and the minimal history length required. We further explore the feasibility of creating a minimized observation policy based on these information. A observation with only $cmd$, $q$, $g$ and $a_{t-1}$ with a window of $T=7$ is equivalent to a mask of 71% on the complete observation space. When directly deployed with such mask, the policy cannot perform well due to all the missing information. As there is no more observation, we freeze the transformer for a fast finetuning on the projection layers and test both the vanilla PPO [27] and supervised learning with online correction [16]. The performance of the policies is shown in Figure 3(d). While both algorithms can help improve the performance of the policy with only minimized observations, supervised learning gives a more significant boost. Training with the teacher has been identified as one major approach to achieve quadruped locomotion on challenging terrains [4, 16]. Although our foundation with the transformer can provide a solid start point of student policy, additional work is still needed for pure RL-based fine-tuning to reduce the dependence on privilege information.
Model Size. We adjust our model by stacking $[1,2,3,4]$ attention blocks, resulting in a model with 0.2M, 0.33M, 0.46M and 0.59M parameters respectively. Figure 3(e) shows that the model performance increases with the model size reaching optimal results with three stacked layers.
@@
Figure 4: Deployment in the physical world with Unitree A1.
5.4 Physical Deployment
@@
We successfully deployed the trained policy, exported with JIT, directly on a Unitree A1 robot equipped with a Jetson AGX Orin Developer Kit. The Jetson acts as both the main processor and a payload. No further model optimization is required for a zero-shot transfer. With the onboard processing power, the policy can process the full observation space at 75Hz and 100Hz while masking out half of the data. With our minimized observation, it can achieve 150Hz, meeting the requirements for real-time deployment and allowing room for further extension with high-dimensional sensors. Figure 4 shows some snapshots from the deployment test. Please refer to the supplementary video for more information.
6 Conclusion
@@
This paper introduces Masked Sensory-Temporal Attention, a transformer-based model for quadruped locomotion. It leverages the masking technique and the direct sensor-level attention to enhance the understanding and generation of sensory information input. We evaluate the robustness of our policy with different combinations of proprioceptive information and demonstrate its capability to compensate for missing data and handle unseen information. Finally, we show that our model is efficient to be deployed on a physical robot without any additional optimization.
@@
Limitation and Future Work. Attention in the full sensory-temporal observation space is computationally intensive and time-consuming. Although using masking can significantly reduce the resources needed, it still takes hours for knowledge transfer, which is considerably longer than existing methods like TCN and temporal-level attention. Additionally, fine-tuning the model with pure reinforcement learning remains challenging, necessitating a more efficient knowledge transfer solution to leverage privileged information effectively. While the policy demonstrated its capability to handle missing data, an addition module is needed to detect and mask out the defected sensors.
Appendix A Quadruped Sensor Set
When selecting sensors for proprioceptive information gathering, we referred to some of the most popular models that are commonly used. Table 2 summarizes the default sensor sets that are typically included out of the box. For low-level tasks such as locomotion, joint encoders and IMUs are the most commonly used sensors, with foot contact sensors also being utilized in some works [4, 21].
Table 2: Sensor set for commonly used quadrupedal robots.
  Robot Model
Joint
Encoder
IMU
Foot
Contact
RGB(-D)
Camera
LiDAR
  A1
Go2
Mini Cheetah
ANYmal C
Spot
Minitaurs
Appendix B Domain Randomization
We provide details of the domain randomization techniques used in simulation in Table 3. To simulate user commands, we frequently resample linear commands in longitudinal and lateral directions, as well as the angular heading, which are then converted to angular velocity commands. To enhance the robustness and generalization of our policy, key robot control factors such as stiffness, damping, and motor strength are randomized. External interferences, such as payload variations and push forces, are also randomized. To accommodate the processing time of the transformer, each action can be delayed by up to 15 ms. The same randomization techniques are applied to both the teacher training session and the knowledge transfer session.
Table 3: Simulation parameters of domain randomization.
  Parameters Range Unit
  Linear Command [-1, 1] $m/s$
Angular Heading [-3.14, 3.14] $rad$
$K_{p}$ Scale [0.9, 1.1] -
$K_{d}$ Scale [0.9, 1.1] -
Friction Scale [0.7, 1.3] -
Motor Strength Scale [0.9. 1.1] -
Payload [0, 5] $kg$
Payload CoM Offset [-0.1, 0.1] $m$
External Push [-1, 1] $m/s$
Gravity [9,41, 10.21] $m/s^{2}$
System Delay [0, 0.015] $s$
Appendix C Reward Terms
Following the approaches in  [21, 4, 26], the rewards are designed to encourage the agent to follow the commanded velocity. To achieve a smooth, safe, and efficient policy, we penalize unwanted motion, sudden changes in actions, feet slippage, body collisions, and excessive power consumption. During testing, we noticed that adding a reward component for balanced power distribution across joints helps to achieve more balanced motion.
Table 4: Reward terms for reinforcement learning
  Reward Definition Scale
  Linear Velocity Tracking $\exp{[-4(v_{xy}^{cmd}-v_{xy})]}$ 1.0
Angular Velocity Tracking $\exp{[-4(\omega_{z}^{cmd}-\omega_{z})]}$ 0.5
Body Z Velocity $\|v_{z}\|^{2}$2.0
Body Rotation $\|\omega_{z}\|^{2}$0.05
Joint Acceleration $\|\ddot{\theta}\|^{2}$2.5e-7
Output Work $\int\|\tau\cdot\dot{\theta}\|$2e-5
Power Distribution $\mathrm{var}(|\tau\cdot\dot{\theta}|)$1e-5
Action Rate $(a_{t}-a_{t-1})^{2}$0.1
Action Smoothness $(a_{t}-2a_{t-1}+a_{t-2})^{2}$0.1
Feet Slip $\|g_{t}\cdot v_{xy}^{feet}\|$0.04
Collision $\mathds{1}_{collision}$1
Appendix D Training Hyper-Parameters
In Table 5, present the detailed training hyper-parameters for PPO training of the teacher policy and the knowledge transfer for transformer-based student policies. The two stages in the knowledge transfer process share the same hyper-parameters.
Table 5: Hyper-Parameters for different training session
  Parameters PPO Supervised Transfer
  Number of GPUs 2 2
Actors per GPU 4096 2048
Episode Length 20s 20s
Horizon Length 24 24
Number of Epochs 5 5
Minibatch Size 16384 3072
Learning Rate 3e-4 5e-4
Scheduler adaptive cosine
Optimizer Adam AdamW
Clip range 0.2 -
Entropy coefficient 0.001 -
Reward Discount 0.99 -
GAE Discount 0.95 -
Desired KL-divergence 0.008 -
Weight Decay - 0.1
Appendix E Extension with Height Information
To better illustrate our framework’s capabilities, we tokenize height map information using a vanilla MLP encoder and add it directly to the observation for fine-tuning over 200 epochs. The normalized test results are shown in Table 6. The inclusion of height information significantly improves the agent’s ability to navigate through terrains, particularly on staircases, which pose the greatest challenge for a small quadruped.
Table 6: Performance with Additional Height Information
  Terrain Metric w/o Height w/ Height
  Smooth Slope Linear Tracking 0.96 0.98
Angular Tracking 0.95 0.96
Total Reward 0.95 0.95
Rough Slope Linear Tracking 0.96 0.96
Angular Tracking 0.91 0.95
Total Reward 0.92 0.93
Stairs Up Linear Tracking 0.89 0.96
Angular Tracking 0.87 0.94
Total Reward 0.76 0.93
Stairs Down Linear Tracking 0.92 0.96
Angular Tracking 0.88 0.94
Total Reward 0.77 0.91
Discrete Linear Tracking 0.92 0.96
Angular Tracking 0.89 0.94
Total Reward 0.83 0.91
Appendix F Learnable Mask
We conduct additional tests to compare the performance of using a learnable representation with the same two-stage transfer. The results, shown in Table 7, indicate that when deployed with masks, the learnable representation cannot match the performance of the removal approach and has a significantly negative impact when all information is present.
Table 7: Performance with Additional Height Information
  Terrain Metric Learnable Ours
0% 25% 50% 0% 25% 50%
  Smooth Slope Linear Tracking 0.21 0.93 0.99 1.00 0.99 0.99
Angular Tracking 0.35 0.97 0.99 1.02 1.02 1.00
Total Reward 0.25 0.92 0.99 1.01 1.01 0.99
Rough Slope Linear Tracking 0.21 0.87 0.95 0.97 0.94 0.95
Angular Tracking 0.33 0.94 0.96 1.01 1.00 0.98
Total Reward 0.19 0.80 0.91 0.98 0.95 0.95
Stairs Up Linear Tracking 0.26 0.81 0.85 0.93 0.91 0.91
Angular Tracking 0.71 0.94 0.93 1.00 0.98 0.97
Total Reward 0.54 0.74 0.72 0.95 0.91 0.87
Stairs Down Linear Tracking 0.24 0.89 0.91 0.94 0.93 0.93
Angular Tracking 0.70 0.93 0.93 1.00 0.99 0.99
Total Reward 0.49 0.75 0.74 0.94 0.91 0.91
Discrete Linear Tracking 0.25 0.87 0.85 0.96 0.96 0.94
Angular Tracking 0.37 0.94 0.94 1.01 1.01 0.99
Total Reward 0.03 0.80 0.76 1.00 0.97 0.96

### v2

@@ V Experiments and Results
We design and conduct various simulation experiments to evaluate the effectiveness of the proposed MSTA, and its generalization ability for different sensor data. We mainly adopt three metrics: linear velocity tracking return per step, angular velocity tracking return per step, and total final reward return. They indicate how the agent can conduct the task following users’ commands and the overall performance. All reported results are averaged over 5000 trails with five terrain types and different levels. They are normalized on the basis of respect teacher data for easy comparison.
V-A Impact of Mask Ratio
@@
First, we investigate the maximum portion of missing data that MSTA can handle to reconstruct robot states. During each the transfer stages, we set the masking ratio to 0%, 25%, 50% and 75% independently. Fig. 3 shows the resultant heatmap matrix. When trained without masking, despite the model having very good performance with all the information available, it suffers from missing data and cannot efficiently reconstruct the status. We can also see that the performance is more dependent on the masking ratio in the second stage than that in the first stage. This is because in the second transfer stage, the student is interacting with the environment to reduce the gap caused by missing information and observation shift. In contrast, the mission of the first stage is to generate a usable policy that outputs reasonable actions so the agent does not fail dramatically and has the chance in the second stage to generate high-quality trajectories for optimization, which is achievable even with a masking ratio of 75%. This also demonstrates the importance of using two-stage transfer.
@@
Comparing the performance of these models, we choose the one trained with the masking ratio of 75% in the first stage and 50% in the second stage, which can well balance the resource requirement and agent performance.
@@
Figure 3: Heatmap matrix for the performance of models that are trained with different combinations of mask ratios. The three rows from top to bottom represent the linear velocity tracking, angular velocity tracking and total reward return respectively. The four columns denote different masking ratios applied during testing. For each sub-figure, the y-axis is the masking ratio applied during the offline pretraining stage and the x-axis is the masking ratio applied during the online correction stage.
@@
TABLE I: Comparison results on different terrain types in terms of linear velocity tracking, angular velocity tracking and total reward return for all the variations of trained models.
  Terrain Metric Ours RMA TERT GRU TERT w/ Mask Ours w/ Learnable
0% 25% 50% Concat Seperate 0% 25% 50% 0% 25% 50% 0% 25% 50%
  Smooth Slope Linear Tracking 1.00 0.99 0.99 1.00 1.01 1.00 1.00 0.96 0.78 0.95 0.98 0.99 0.21 0.93 0.99
Angular Tracking 1.02 1.02 1.00 1.02 1.02 1.02 1.02 1.01 0.97 0.99 1.00 0.99 0.35 0.97 0.99
Total Reward 1.01 1.01 0.99 1.02 1.02 1.02 1.01 0.99 0.84 0.95 0.99 0.98 0.25 0.92 0.99
Rough Slope Linear Tracking 0.97 0.94 0.95 0.99 1.00 0.98 0.96 0.90 0.74 0.91 0.95 0.93 0.21 0.87 0.95
Angular Tracking 1.01 1.00 0.98 1.02 1.02 1.01 1.01 1.00 0.95 0.96 0.97 0.94 0.33 0.94 0.96
Total Reward 0.98 0.95 0.95 1.00 1.01 0.99 0.97 0.92 0.79 0.90 0.94 0.89 0.19 0.80 0.91
Stairs Up Linear Tracking 0.93 0.91 0.91 0.94 0.95 0.95 0.90 0.91 0.69 0.82 0.86 0.85 0.26 0.81 0.85
Angular Tracking 1.00 0.98 0.97 0.99 1.00 1.01 0.99 0.98 0.94 0.95 0.95 0.92 0.71 0.94 0.93
Total Reward 0.95 0.91 0.87 0.95 0.99 1.00 0.90 0.87 0.72 0.80 0.81 0.73 0.54 0.74 0.72
Stairs Down Linear Tracking 0.94 0.93 0.93 0.96 0.97 0.95 0.92 0.93 0.74 0.86 0.91 0.92 0.71 0.94 0.93
Angular Tracking 1.00 0.99 0.99 1.00 1.01 1.01 0.99 0.98 0.95 0.95 0.96 0.92 0.70 0.93 0.93
Total Reward 0.94 0.91 0.91 0.95 1.00 0.98 0.89 0.90 0.77 0.83 0.86 0.77 0.49 0.75 0.74
Discrete Linear Tracking 0.96 0.96 0.94 0.97 0.99 0.98 0.93 0.88 0.75 0.87 0.91 0.89 0.25 0.87 0.85
Angular Tracking 1.01 1.01 0.99 1.01 1.02 1.02 1.01 0.99 0.95 0.96 0.97 0.94 0.37 0.94 0.94
Total Reward 1.00 0.97 0.96 0.99 1.04 1.02 0.94 0.89 0.83 0.88 0.91 0.80 0.03 0.80 0.76
  Average Linear Tracking 0.96 0.95 0.94 0.97 0.98 0.97 0.94 0.92 0.74 0.88 0.92 0.92 0.23 0.87 0.91
Angular Tracking 1.01 1.00 0.99 1.01 1.01 1.01 1.00 0.99 0.95 0.96 0.97 0.94 0.49 0.95 0.95
Total Reward 0.97 0.95 0.94 0.98 1.01 1.00 0.94 0.92 0.79 0.87 0.90 0.83 0.30 0.80 0.82
V-B Comparison with Baselines
We compare MSTA with two baselines. The first is RMA [4], which is implemented with TCN [34] to capture temporal information. The second is TERT [15], a transformer-based framework with linear projection for observations and actions in two favors: concatenated single token and separate tokens for states and action, resulting in $T$ and $2T$ tokens respectively [17]. To evaluate the masking mechanism in our method, we replace the selected observation in MSTA with a learnable representation instead of removing them. To further evaluate the importance and capability of the transformer structure, we replace it with a GRU [35] model. We expand the missing information testing to TERT. However, since the observations and actions in TERT are encoded through linear projection before passing to the transformer, it is impossible to directly remove any input. Thus, the the same learnable masks method is applied to TERT.
All variations of MSTA, TERT and other baselines are trained with the same two-stage transfer, sharing a common well-trained teacher network, and we apply a testing mask of up to 50%, as identified in Section V-A.
Tab. I shows the comparison results. When fully optimized with teacher-student transfer, the performance of all fully trained vanilla policies with complete observations is very close, often within just 2% difference. When faced with incomplete information, transformer-based MSTA can have a better understanding of the data and reconstruct the robot state more accurately than the GRU-based network, even with only half of the information. When using a learnable representation mask, with MSTA or TERT, the agent underperforms to the vanilla removing mask, especially with full observation, showing that a direct removing mask has an advance in both better performance and less resource required. Although we can hack the linear projection in the TERT network to take in missing information, it is not comparable to direct sensor level tokenzation and attention for sensory information understanding.
V-C Generalization, Robustness and Flexibility
While achieving state-of-the-art performance, MSTA offers additional benefits of generalization and flexibility to customize the model after training or even on the fly to fit the deployment requirement. Quadrupeds are equipped with different sensor sets, and sensor damage can cause certain channels or the entire sensor to be unavailable during deployment, which required the robustness against missing information to handle. Furthermore, we can balance the performance and required computation power by using a shorter sequence based on the insights from in-context sensory information understanding.
Figure 4: Performance with certain sensory feedback completely removed.
@@ V-C Generalization, Robustness and Flexibility
Figure 5: Performance with various setups: Left certain numbers of joint encoders are masked out; Right different history time window $T$ is applied.
@@
Important Sensory Feedback. To understand the importance of each sensory feedback, we further investigate the impact of removing each sensor completely from the observation and the results are shown in Fig. 4. It is clear that certain feedback like $\dot{q}$, $c$, $\omega$ and even $a_{t-1}$ are quite redundant and a well trained transformer-based MSTA can easily compensate the missing information from other sources, while the other sensor data are more critical.
@@
Missing of Sensor Dimension. Some proprioceptive information has multiple channels, such as joint encoders and force senors. This means that these sensors can also be damaged independently due to wear and tear from daily operations and it is not easy to have a redundant sensor. Among these sensors, joint encoders are the source of both $q$ and $\dot{q}$ for the observation. From previous analysis, missing of joint information can be crucial. We investigate the scenario where only a few encoders are dead or the data are compromised and need to be excluded. We conduct the test by masking certain numbers of joint encoders and for each masked joint, the related $q$ and $\dot{q}$ are removed completely from the observation. The results are shown in Fig. 5. The loss of the joint encoder can have a great impact on the performance as the related information is very essential for quadruped locomotion. However, our transformer model can still handle multiple missing encoders before large performance degradation.
@@
Time Window. Another special masking is to completely remove some timesteps, the default window, $T=15$, is equivalent to past 0.3s. We check whether such a long sequence of information is necessary by applying different time windows without masking. The results are shown in Fig. 5. It is clear that MSTA can efficiently extract and reconstruct the robot state for actions even with only 7 steps of past information. Interestingly, given a longer timeframe like $T=20$, which the transformer has never seen during training, MSTA is still robust and not affected by such unknown information.
Figure 6: Performance using minimized observations with finetuning and extension of height map.
Minimized Observation and Fine-tuning. It is not feasible to infer the transformer with full observation space and long windows with the limited computation power onboard. From the previous analysis, we have identified the important sensors and the minimal history length required. We further explore the feasibility of creating a minimized observation policy based on the information. Using an observation with only $cmd$, $q$, $g$ and $a_{t-1}$ with a window of $T=7$, we are essentially removing 71% of the tokens from the complete training observation space.
@@
When directly deployed with such mask, the policy cannot perform well due to all the missing information. To restore the performance, we freeze the transformer for fast fine-tuning of the projection layers and test both the vanilla PPO [29] and supervised learning with online correction [15]. The performance of the policies is shown in Fig. 6. While both algorithms can help improve the performance of the policy with only minimized observations, supervised learning gives larger boost. Training with the teacher has been identified as one major approach to achieve quadruped locomotion on challenging terrains [4, 15]. Although our foundation with the transformer can provide a solid start point of student policy, additional work is still needed for pure RL-based fine-tuning to reduce the dependence on privilege information.
Extension with New Information. In previous analysis, MSTA is robustness against new timestep information. When extending the capability for quadrupeds, additional sensors such as cameras and LiDAR are often needed. We assess the model’s capability of handling previously unseen information, which can be appended into $\mathcal{T}$ as new tokens. For instance, we tokenize the height map information using a vanilla MLP encoder and directly extend it to our minimized observation agent for fine-tuning. The performance of the extended agent is shown in Fig. 6. Height information significantly aids in navigating challenging terrains, such as staircases, and improves the overall locomotion performance even with minimal observations and new encoder needed to be trained. To take the test to an extreme, we added 256 randomly generated dummy tokens, equivalent to a camera frame with ViT [33] before processing, and the agent can still produce explore, which is crucial for two-stage knowledge transfer. This demonstrates that the model can be used as a solid foundation for further extension with high-dimensional information by direct deployment in virtual environments to gather new trajectories. Please refer to the supplementary video for more information.
@@
Figure 7: Deployment in the physical world on Unitree A1 with minimized observations with zero-shot transfer.
V-D Physical Deployment
@@
We successfully deploy the trained policy, exported with JIT, directly on a Unitree A1 robot equipped with a Jetson AGX Orin Developer Kit. The Jetson acts as both the main processor and a payload. No further model optimization is required for a zero-shot transfer. With the onboard processing power, the policy can run at 150Hz with our minimized observation, meeting the requirements for real-time deployment and allowing room for further extension with high-dimensional sensors. However, we notice that the JIT model will have slight difference in output compared to the original model, indicating additional work is needed for better portability. Fig. 7 shows some snapshots from the deployment test. Please refer to the supplementary video for more information.
VI Conclusion
@@
This paper introduces MSTA, a transformer-based model for quadruped locomotion. It leverages the masking technique and direct sensor-level attention to enhance the understanding and generation of sensory information input. We evaluate the robustness of MSTA with different combinations of proprioceptive information and demonstrate its capability to compensate for missing data and handle unseen information. Finally, we show that MSTA is efficient to be deployed on a physical robot without any additional optimization.
@@
Attention in the full sensory-temporal observation space is computationally intensive and time-consuming. Although using masking can significantly reduce the resources needed, it still takes hours for knowledge transfer, which is considerably longer than existing methods like TCN and temporal-level attention. Additionally, fine-tuning the model with pure reinforcement learning remains challenging, necessitating a more efficient knowledge transfer solution to leverage privileged information effectively. While the policy demonstrated its capability to handle missing data, an addition module is needed to detect and mask out the defected sensors. These will be our future work.